OpenStudy (anonymous):

Which rule can you use to find the nth term of an arithmetic sequence in which the common difference is 5 and a12 = 63?

OpenStudy (anonymous):

A an=3+5 B an=3-5n C an=13+5n D an=-5n im thinking its c but im not sure?

OpenStudy (anonymous):

let a be the first term tn=a+(n-1)d when n=12,tn=63 63=a+(12-1)*5 a=63-55=8 tn or an=8+(n-1)5=8+5n-5=3+5n

OpenStudy (anonymous):

so thatd be A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank u so much!

OpenStudy (anonymous):

you have written only 3+5 typing error.